A number cube is rolled and a coin is flipped. If the number cube shows and odd number, or the coin shows tails, or both, you get 6 points. If neither event occurs, you get 4 points. What is the expected value of this game?
Answer by edjones(8007)

1/2 * 1/2 = 1/4 of the time the number of the die will be even and the coin will show heads and you get 4 pts.
3/4 of the time you get 6 pts.
3/4 * 6 + 1/4 * 4
=9/2 + 1
=11/2 = 5.5 points, the expected value of the game.
.
Ed